# Pushing the Envelope of Low-Bit LLM via Dynamic Error Compensation

### v1

@@ 4.4 Parameter Tuner
Necessity of Parameter Tuner. To effectively utilize QDEC, two key parameters must be carefully tuned. The first is the ratio of input channels to compensate in each layer, controlled by $k_{chunk}$, which denotes the number of input channels to compensate per 1024 input channels. A larger $k_{chunk}$ improves model quality but may increase inference latency. The second is $n_{tb}$, the number of thread blocks allocated for the dynamic error compensation. Since the dynamic error compensation runs in parallel with the base GEMV, assigning too many thread blocks to it can slow down the base GEMV. On the other hand, if too few thread blocks are assigned, the PCIe bandwidth cannot be fully utilized, as the zero-copy mechanism relies on GPU cores to issue memory requests.
@@
Determining $k_{chunk}$ and $n_{tb}$ requires significant effort due to the vast design space. The value of $k_{chunk}$ for each type of linear layer in a decoder block ($k^{qkv}_{chunk}$, $k^{o}_{chunk}$, $k^{gu}_{chunk}$, $k^{d}_{chunk}$) can be any integer less than its maximum possible value, which is constrained by the shared memory size. While this maximum may vary across platforms, it is typically quite large (e.g., 367 in the NVIDIA Ada architecture), resulting in an expansive search space. Meanwhile, $n_{tb}$ values for each layer ($n^{qkv}_{tb}$, $n^{o}_{tb}$, $n^{gu}_{tb}$, $n^{d}_{tb}$) are constrained by the input and output dimensions of the weights, leading to far fewer possible candidates compared to $k_{chunk}$. Still, however, each layer has multiple viable candidate values. For example, in Llama-3-8B, there are 9 possible candidates for $n^{qkv}_{tb}$ (1, 2, 3, 4, 5, 6, 8, 12, 24), while other layers also have multiple options. This leads to a combinatorial explosion in the number of possible configurations. We provide details about identifying possible values for $k_{chunk}$ and $n_{tb}$ in the supplementary material.
@@
To ease this process, we provide a parameter tuner for QDEC, which suggests values for $k_{chunk}$ and $n_{tb}$ based on the user’s target slowdown rate. Specifically, the tuner seeks to maximize $k_{chunk}$ values while ensuring that the total execution time—including the base GEMV and the dynamic error compensation across all linear layers in the block—remains within the target slowdown rate compared to the baseline execution time without dynamic error compensation. This parameter tuning is a one-time process for a given model-device pair.
@@
Paramter Tuning Process. Figure 11(a) illustrates the tuning process, which consists of two phases: Phase 1 determines the $n_{tb}$ values, and Phase 2 determines the $k_{chunk}$ values.
@@
In Phase 1, the tuner simplifies the search for $n_{tb}$ values for each layer by replacing it with the search for a single metaparameter, $n^{max}_{tb}$. $n^{max}_{tb}$ represents the upper limit on the number of thread blocks to use for dynamic error compensation. Once the $n^{max}_{tb}$ value is determined, the $n_{tb}$ values for each layer are set to the largest possible candidate value smaller than $n^{max}_{tb}$. Determining $n^{max}_{tb}$ involves testing values up to half of the total SM count, rather than up to the total SM count, to reduce the search space. In the example shown in Figure 11, the GPU has 56 SMs, so values up to 28 are tested. Each test includes a coarse-grained search for $k_{chunk}$, which is to determine, for the given $n^{max}_{tb}$, how many steps can be taken when incrementing $k_{chunk}$ uniformly for all layers without exceeding the target slowdown rate. Figure 11(b) shows an example of a coarse-grained $k_{chunk}$ search for $n^{max}_{tb}=24$ (i.e., $(n^{qkv}_{tb},n^{o}_{tb},n^{gu}_{tb},n^{d}_{tb})=(24,16,23,14)$), where a total of 19 steps were made. If no steps can be made for any $n^{max}_{tb}$ value, the tuner fixes $k_{chunk}$ to 0 for the layer with the smallest weight matrix and repeats the process, as smaller matrices are often most sensitive to increases in $k_{chunk}$.
@@
After Phase 1, the tuner selects the $n^{max}_{tb}$ that allows the most steps and proceeds to Phase 2, which involves a fine-grained $k_{chunk}$ search. Figure 11(c) illustrates this fine-grained $k$ search with the selected $n^{max}_{tb}$, which is 24 in this example. In this phase, not all $k_{chunk}$ values may increase together. At each step, the tuner increments $k_{chunk}$ for as many layers as possible, prioritizing those with smaller increases in execution time. For example, in Step 1, $k^{gu}_{chunk}$ is incremented first, followed by $k^{d}_{chunk}$. Step 1 stops at this point, as further increases to $k^{qkv}_{chunk}$ and $k^{o}_{chunk}$ would exceed the target; thus their final values are set (i.e., $k^{qkv}_{chunk},k^{o}_{chunk}=19$). This process repeats until no further increments can be made for any layer.
@@ 5 Evaluation
5.1 Methodology
Quantized LLMs. We evaluate the effectiveness of QDEC by integrating it on 3-bit, 3.5-bit and 4-bit versions of two instruction-tuned LLMs: Llama-3-8B-Instruct [16] and Phi-3-medium-4k-instruct (14B) [39]. For the 3.5-bit version, we adopt a block-wise bitwidth allocation, applying 3-bit quantization to half of the decoder blocks and 4-bit quantization to the remaining blocks. This allocation is guided by a Kullback–Leibler (KL) divergence-based sensitivity metric, as proposed in prior work [8]. As base quantization methods, we choose two state-of-the-art LLM quantization methods that excel in both model quality and inference latency: AWQ [35] and SqueezeLLM [30]. AWQ is a uniform quantization method that mitigates quantization errors by applying per-channel scaling to protect salient channels, which are identified through an offline analysis on a calibration dataset. SqueezeLLM employs a clustering-based non-uniform quantization method that considers the sensitivity of each weight.
Benchmarks. Following previous literature [9, 30, 35, 12], we use perplexity on the WikiText dataset [38] as a primary metric since perplexity reliably reflects the quality of quantized LLMs [35, 13]. Additionally, we use BIG-Bench Hard (BBH) [53], a collection of 23 challenging tasks in BIG-Bench [51], to assess the models’ capabilities in complex problem solving. Chain-of-Thought (CoT) prompting is enabled for all BBH evaluations [61]. Lastly, we evaluate multi-turn conversation using MT-Bench [67], where a strong LLM serves as a judge, scoring the model’s responses on 80 sample questions with integer scores ranging from 0 to 10. We use GPT-4o as the judge. For MT-Bench evaluation, we repeat the experiments three times and report the average score.
Figure 12: Perplexity on WikiText. The x markers correspond to baselines without QDEC ($k_{chunk}=0$). Lower is better.
Figure 13: Accuracy on BBH. The x markers correspond to baselines without QDEC ($k_{chunk}=0$). Higher is better.
Figure 14: MT-Bench scores. The x markers correspond to baselines without QDEC ($k_{chunk}=0$). Higher is better.
@@ 5.2 Impact on Model Quality
Methodology. In this section, we demonstrate the model quality improvements achieved by QDEC. We evaluate models with varying $k_{chunk}$, the number of channels compensated per chunk (1024 channels), uniformly setting it to 8, 16, 32, 64 and 128 for all layers.
@@
Perplexity on WikiText. Figure 12 shows the perplexity results. Across all cases, a clear trend is observed: perplexity consistently decreases (i.e., model quality improves) as $k_{chunk}$ increases. Notably. for 3-bit models, significant drops in perplexity are achieved even with a low $k_{chunk}$ of 8. AWQ’s perplexity decreases from 10.15 to 9.63 for Llama-3 and from 5.96 to 5.53 for Phi-3, while SqueezeLLM’s perplexity drops from 10.49 to 9.93 for Llama-3 and from 5.92 to 5.45 for Phi-3. Meanwhile, the impact of dynamic error compensation on 4-bit models is relatively less pronounced. This is expected, as 4-bit models already offer high quality and are much closer to the full-precision models, leaving a smaller room for improvement. The 3.5-bit models exhibit trends that fall between those of the 3-bit and 4-bit models.
@@
BBH and MT-Bench. Results on BBH (Figure 13) follow the same trends as the perplexity results. The MT-Bench results (Figure 14), on the other hand, demonstrate some different patterns. In cases where vanilla quantized models without QDEC ($k=0$) already achieve scores very close to those of the FP16 model—such as all 4-bit cases and the AWQ 3.5-bit model of Phi-3—the scores remain unchanged with QDEC, oscillating around the baseline score. For the remaining cases, QDEC significantly improves scores even with a small $k_{chunk}$ (e.g., 8), like in other benchmarks; further increases in $k_{chunk}$, however, do not always yield noticeable improvements. These patterns may be attributed to the coarse-grained rubric of this benchmark, which assigns integer scores ranging from 0 to 10 for each task. Such a scoring scheme may fail to capture the subtle improvements introduced by dynamic error compensation when the potential for improvement is limited.
Figure 15: Comparison against exact Top-K channel selection. line graphs show perplexity, and bar graphs represent the average recall rate of QDEC’s approximate Top-K.
Impact of Top-K Approximation. Figure 15 compares the perplexity between QDEC and QDEC with exact Top-K selection to investigate the impact of QDEC’s Top-K approximation. The average recall rate of QDEC’s approximate Top-K is also reported. In all cases, the approximation performs well, achieving a recall rate of around 80% and yielding perplexity nearly identical to the exact Top-K.
Figure 16: Comparison against static and random channel selection. Lower is better.
Impact of Dynamic Salient Channel Selection. To evaluate the impact of dynamic salient channel selection in QDEC, we compare its perplexity with its two variants: (1) compensating for errors in a static set of channels identified via calibration set analysis, and (2) compensating for errors in a random set of channels. For the static identification of salient channels, we use a Hessian-based metric following prior work [32]. Figure 16 shows the results. While QDEC with static channel selection outperforms random selection, it significantly falls short of the original dynamic QDEC. For example, to achieve a similar level of perplexity as QDEC with $k_{chunk}=64$, QDEC with static channel selection requires $k_{chunk}$ to be 256, making it impractical.
@@
Figure 17: Execution time of base GEMV + QDEC with varying $k_{chunk}$ and $n_{tb}$, normalized to base GEMV execution time.
5.3 GPU Kernel Benchmarks
GPU Name Memory Size Memory BW # SM PCIe BW
Desktop RTX 4090 24 GB 1,008 GB/s 128 32 GB/s
RTX 4080S 16 GB 736 GB/s 80 32 GB/s
RTX 4070S 12 GB 504 GB/s 56 32 GB/s
Laptop RTX 4070M 8 GB 256 GB/s 36 16 GB/s
RTX 4050M 6 GB 192 GB/s 20 16 GB/s
Table 2: GPU specifications
Methodology. In this section, we evaluate the QDEC GPU kernel on three consumer-grade GPUs: two desktop GPUs, RTX 4070 Super (RTX 4070S) and RTX 4090, and one laptop GPU, RTX 4050 Mobile (RTX 4050M), with their specifications detailed in Table 2. The evaluation considers GEMV operations of output, gate/up and down projection layers in Llama-3, assuming a 3-bit bitwidth. We use the GEMV kernel from LUTGEMM [44] for AWQ and the Any-Precision LLM kernel [45] for SqueezeLLM, both state-of-the-art GEMV kernels for uniform and non-uniform quantization, respectively. Kernel times are measured using NVIDIA Nsight Systems.
Results. Figure 17 shows the execution time of QDEC kernel (base GEMV + dynamic error compensation), normalized to the standalone execution time of the base GEMV, across varying $n_{tb}$ and $k_{chunk}$. The expected behavior of QDEC’s kernel time resembles a piecewise linear function with two segments: the first segment remains constant near the base GEMV time (i.e., normalized latency $\sim$ 1.0), corresponding to the region where dynamic error compensation operations are hidden under the base GEMV; the second segment increases linearly with $k_{chunk}$, representing the region where these operations are no longer hidden. The knee point, where the transition occurs, is preferable at larger values of $k$.
All cases exhibit the expected behavior when $n_{tb}$ is properly set, except for the $4096\times 4096$ case on RTX 4090, the fastest GPU evaluated. In this case, the base GEMV execution time is so short that even a small $k_{chunk}$ incurs latency overhead. From the other cases, three key observations can be made. First, the knee point is highly sensitive to the choice of $n_{tb}$, highlighting the importance of careful tuning. Generally, higher $n_{tb}$ values such as 8 or 16 delay the knee point. In contrast, smaller $n_{tb}$ values like 2 lead to suboptimal performance by causing the knee point to occur too early or not at all. However, increasing $n_{tb}$ can sometimes worsen results by slowing down the base GEMV, as discussed in Section 4.4. This effect is particularly noticeable on GPUs with fewer SMs, such as the 4050M. For instance, on the 4050M, $n_{tb}$ values of 4 or 8 seem to yield the best results, whereas increasing $n_{tb}$ to 16 leads to worse performance. Second, large weight sizes allow for a larger $k_{chunk}$ with less latency increase, as they provide more time slack. Third, the lower the GPU memory bandwidth and the higher the PCIe bandwidth, the later the knee point. Thus, RTX 4050M, with the highest PCIe-to-GPU memory bandwidth ratio, allows the largest $k_{chunk}$ before the knee point; RTX 4090, with the lowest ratio, allows the smallest.

### v2

@@ 4.4 Parameter Tuner
Necessity of Parameter Tuner. Effective use of DecDEC requires careful tuning two key parameters:
  • $n_{tb}$: Specifies the number of thread blocks used for dynamic error compensation. Since DecDEC runs in parallel with the base GEMV, allocating too many thread blocks to dynamic error compensation can slow down the base GEMV computation, while allocating too few may underutilize PCIe bandwidth, as zero-copy transfers require GPU cores to issue memory requests.
  • $k_{chunk}$: Specifies how many channels are compensated per chunk (1024 channels). A larger $k_{chunk}$ improves model quality but may increase inference latency.
@@
Choosing $n_{tb}$ and $k_{chunk}$ is challenging due to the large design space. The $n_{tb}$ value for each type of linear layer ($n^{qkv}_{tb}$, $n^{o}_{tb}$, $n^{gu}_{tb}$, $n^{d}_{tb}$) has multiple viable candidates, constrained by the dimensions of the corresponding weight matrices. For example, in Llama-3-8B, there are 9 possible candidates for $n^{qkv}_{tb}$ (1, 2, 3, 4, 5, 6, 8, 12, 24), while other layers also have multiple options. Selecting $k_{chunk}$ presents a greater challenge due to its broader range of possibilities. The $k_{chunk}$ value for each type of linear layer ($k^{qkv}_{chunk}$, $k^{o}_{chunk}$, $k^{gu}_{chunk}$, $k^{d}_{chunk}$) can be any integer less than a maximum determined by the available shared memory. Although platform-dependent, this upper bound is large (e.g., 367 with 48 KB per-block shared memory), leading to an expansive search space. This creates a combinatorial explosion of configuration options. Technical details on how to identify candidate values of $k_{chunk}$ and $n_{tb}$ for a given model and platform are presented at the end of this section.
@@
To address this issue, we provide a DecDEC tuner that suggests $n_{tb}$ and $k_{chunk}$ values based on a target slowdown rate. The tuner maximizes $k_{chunk}$ while keeping the total execution time—including base GEMV and dynamic error compensation across all linear layers—within the specified slowdown relative to the baseline (i.e., without compensation). This tuning is a one-time process for a given model-device pair.
@@
Parameter Tuning Process. Figure 11(a) illustrates the tuning process, which consists of two phases: Phase 1 determines the $n_{tb}$ values, and Phase 2 determines the $k_{chunk}$ values.
@@
In Phase 1, the tuner simplifies the search for $n_{tb}$ values for each layer by replacing it with the search for a single metaparameter, $n^{max}_{tb}$. $n^{max}_{tb}$ represents the upper limit on the number of thread blocks to use for dynamic error compensation. Each layer’s $n_{tb}$ is then set to the largest candidate below $n^{max}_{tb}$. Determining $n^{max}_{tb}$ involves testing values up to half of the total SM count, to reduce the search space. In the example shown in Figure 11, the GPU has 56 SMs, so values up to 28 are tested. For each tested $n^{max}_{tb}$, a coarse-grained $k_{chunk}$ search checks how many uniform increments to $k_{chunk}$ can be applied across all layers without exceeding the target slowdown rate. Figure 11(b) shows an example of a coarse-grained $k_{chunk}$ search for $n^{max}_{tb}=24$ (i.e., $(n^{qkv}_{tb},n^{o}_{tb},n^{gu}_{tb},n^{d}_{tb})=(24,16,23,14)$), yielding 19 valid steps. If no steps can be made for any $n^{max}_{tb}$ value, the tuner fixes $k_{chunk}$ to 0 for the layer with the smallest weight matrix and repeats the process, as smaller matrices are often most sensitive to increases in $k_{chunk}$.
@@
After Phase 1, the tuner selects the $n^{max}_{tb}$ with the most steps and proceeds to a fine-grained $k_{chunk}$ search in Phase 2. Figure 11(c) shows the fine-grained $k$ search for the selected $n^{max}_{tb}=24$. In this phase, not all $k_{chunk}$ values may increase together. At each step, the tuner increments $k_{chunk}$ for as many layers as possible, prioritizing those with smaller increases in execution time. For example, in Step 1, $k^{gu}_{chunk}$ is incremented first, followed by $k^{d}_{chunk}$. Step 1 stops at this point, as further increases to $k^{qkv}_{chunk}$ and $k^{o}_{chunk}$ would exceed the target; thus their final values are set (i.e., $k^{qkv}_{chunk},k^{o}_{chunk}=19$). This process repeats until no further increments can be made for any layer.
Technical Details: $n_{tb}$ and $k_{chunk}$ Candidates. The $n_{tb}$ values considered during parameter tuning are those that have a meaningful impact on at least one of the two parts of the kernel execution: the approximate Top-K selection and the residual fetching. In the approximate Top-K selection part, the minimum processing granularity per thread block is one chunk. Therefore, increasing $n_{tb}$ beyond the number of chunks does not have additional effect on performance. Thus, the set of $n_{tb}$ values relevant to this part is:
[…]
For residual fetching part, 4-bit residuals are transferred over PCIe in coalesced segments of 256 values (128 bytes), resulting in a total of $s=d_{out}/256$ segments. These segments are distributed across $n_{tb}$ thread blocks, with each block processing $\lceil s/n_{tb}\rceil$ segments. If multiple $n_{tb}$ values result in the same number of segments per block ($\lceil s/n_{tb}\rceil$), only the smallest such value is considered and the rest are redundant. Excluding these cases, the candidate set relevant to this part is:
[…]
The final candidate set for $n_{tb}$ is the union of the two (i.e., $N=A\cup B$).
Meanwhile, $k_{chunk}$ is bounded by the shared memory limit. During the approximate Top-K selection part of the kernel, shared memory usage increases with the value of $k_{chunk}$. Specifically, the shared memory usage for this part is:
$128+128\times k_{chunk}+2\times 1024\;\text{bytes}.$
Here, 128 bytes are used for integer counters that track the number of elements that fall into each of the 32 buckets; $128\times k_{chunk}$ bytes are used to temporarily store the indices of elements assigned to each bucket; and $2\times 1024$ bytes account for the input activation values in the chunk. The total must remain below the per-block shared memory limit (e.g., 49,152 bytes), which constrains the maximum allowable $k_{chunk}$ value.
@@ 5 Evaluation
Figure 12: Execution time of base GEMV + DecDEC with varying $k_{chunk}$ and $n_{tb}$, normalized to base GEMV execution time.
5.1 GPU Kernel Benchmarks
GPU Name Memory Size Memory BW # SM PCIe BW $\bm{R_{bw}}$
Desktop
RTX 4090 24 GB 1,008 GB/s 128 32 GB/s 32
RTX 4080S 16 GB 736 GB/s 80 32 GB/s 23
RTX 4070S 12 GB 504 GB/s 56 32 GB/s 16
Laptop
RTX 4070M 8 GB 256 GB/s 36 16 GB/s 16
RTX 4050M 6 GB 192 GB/s 20 16 GB/s 12
Table 1: GPU specifications.
Methodology. In this section, we evaluate the DecDEC GPU kernel on three consumer-grade GPUs: two desktop GPUs, RTX 4070 Super (RTX 4070S) and RTX 4090, and one laptop GPU, RTX 4050 Mobile (RTX 4050M), with their specifications detailed in Table 1. Here, $R_{bw}$ denotes the ratio of memory bandwidth to CPU-to-GPU (PCIe) bandwidth. The evaluation considers GEMV operations of output, gate/up and down projection layers in Llama-3-8B-Instruct [16], assuming 3-bit bitwidth. For the base GEMV, we use the LUTGEMM kernel,  [44], a state-of-the-art GEMV kernel for uniform quantization. Kernel times are measured using NVIDIA Nsight Systems.
Expected Behavior. The execution time of the DecDEC kernel is expected to follow a piecewise linear function of $k_{chunk}$ with two distinct segments. In the first segment, corresponindg to small $k_{chunk}$ values, the latency remains nearly constant at the base GEMV time, as the compensation operations are fully hidden under the GEMV execution. In the second segment, once $k_{chunk}$ exceeds a certain threshold—the knee point—the execution time linearly increases with $k_{chunk}$, inidcating that the compensation operations are no longer fully hidden. Theoretically, the knee point is expected to occur when $k_{chunk}=1024\times\nicefrac{{1}}{{\text{{$R_{bw}$}}}}\times\nicefrac{{3}}{{4}}$, representing the maximum amount of data transfer that can overlap with the base GEMV. Here, the base GEMV time is approximated by dividing the weight matrix size by the GPU memory bandwidth, and PCIe bandwidth is assumed to be fully utilized. The factor $\nicefrac{{3}}{{4}}$ accounts for the 3-bit quantization; for other bitwidths, this factor should be adjusted accordingly (e.g., $\nicefrac{{4}}{{4}}=1$ for 4-bit quantization).
Results. Figure 12 shows the execution time of DecDEC kernel (base GEMV + dynamic error compensation), normalized to the standalone execution time of the base GEMV, across varying $n_{tb}$ and $k_{chunk}$. Each subfigure also includes a vertical dotted line to mark the theoretical knee point derived from the analytical model (i.e., $1024\times\nicefrac{{1}}{{\text{{$R_{bw}$}}}}\times\nicefrac{{3}}{{4}}$). All cases exhibit the expected two-segment piecewise linear behavior when $n_{tb}$ is properly set, except for the $4096\times 4096$ case on RTX 4090, the fastest GPU evaluated. In this case, the base GEMV execution time is so short that even a small $k_{chunk}$ incurs overhead.
From the other cases, three key observations can be made. First, a lower $R_{bw}$—indicating lower memory bandwidth and higher PCIe bandwidth—shifts the knee point to the right. Thus, RTX 4050M, with the lowest $R_{bw}$, supports the largest $k_{chunk}$ before the knee point, while RTX 4090, with the highest ratio, supports the smallest. This trend is consistent with the ordering predicted by theoretical knee points. Second, the knee point is highly sensitive to $n_{tb}$, highlighting the importance of careful tuning. Generally, higher $n_{tb}$ values such as 8 or 16 delay the knee point. In contrast, small $n_{tb}$ values (e.g., 2) cause the knee point to appear too early or disappear, leading to suboptimal performance. However, increasing $n_{tb}$ can sometimes worsen results by slowing down the base GEMV, as discussed in Section 4.4. This effect is particularly noticeable on GPUs with fewer SMs, such as the 4050M. For instance, on the 4050M, $n_{tb}=8$ yields the best results, whereas increasing $n_{tb}$ to 16 leads to worse performance. Third, larger weight sizes enable higher $k_{chunk}$ with minimal latency overhead, due to increased time slack. With sufficiently large weight sizes (e.g., 4096$\times$28672) and a properly tuned $n_{tb}$, the actual knee point approaches the theoretical value. For example, on an RTX 4050M with a 4096$\times$28672 matrix and $n_{tb}=8$, the observed knee point is around $k_{chunk}=60$, compared to a theoretical prediction of 64.
@@ 5.2 Impact on Model Quality
Methodology. We demonstrate DecDEC ’s quality improvements on 3-bit, 3.5-bit, and 4-bit versions of two instruction-tuned LLMs: Llama-3-8B-Instruct [16] (hereafter referre to as Llama-3) and Phi-3-medium-4k-instruct (14B model, hereafter referred to as Phi-3) [39]. For the 3.5-bit version, we adopt a block-wise bitwidth allocation, applying 3-bit quantization to half of the decoder blocks and 4-bit quantization to the remaining blocks. This follows a KL divergence-based sensitivity metric from prior work [8].
As base quantization methods, we choose two state-of-the-art LLM quantization methods: AWQ [35] and SqueezeLLM [30]. AWQ is a uniform quantization method that mitigates quantization errors by applying per-channel scaling to protect salient channels, which are identified through an offline analysis on a calibration dataset. SqueezeLLM employs a clustering-based non-uniform quantization method that considers the sensitivity of each weight.
We evaluate DecDEC-augmented models across varying $k_{chunk}$, the number of channels compensated per chunk (1024 channels). We uniformly set $k_{chunk}$ to 8, 16, 32, 64 and 128 for all layers. Based on the results in Section 5.1, values up to 64 fall within the practical range—where latency overhead from dynamic error compensation may remain low depending on the platforms—while $k_{\text{chunk}}=128$ is included for completeness in evaluating upper-bound behavior.
Benchmarks. Following previous literature [9, 30, 35, 12], we use perplexity on WikiText [38] as the primary metric, as it reliably reflects quantized LLM quality [35, 13]. Additionally, we use BIG-Bench Hard (BBH) [54], a collection of 23 challenging tasks in BIG-Bench [52], to assess the models’ capabilities in problem solving. Chain-of-Thought (CoT) is enabled for all BBH evaluations [62]. Lastly, we evaluate multi-turn conversation using MT-Bench [68], where a strong LLM judges 80 responses, scoring each from 0 to 10. We use GPT-4o as the judge and report the average of three runs.
Figure 13: Perplexity on WikiText. The x markers correspond to baselines without DecDEC ($k_{chunk}=0$). Lower is better.
Figure 14: Accuracy on BBH. The x markers correspond to baselines without DecDEC ($k_{chunk}=0$). Higher is better.
Figure 15: MT-Bench scores. The x markers correspond to baselines without DecDEC ($k_{chunk}=0$). Higher is better.
@@
Perplexity on WikiText. Figure 13 shows the perplexity results. Across all cases, a clear trend is observed: perplexity consistently decreases (i.e., model quality improves) as $k_{chunk}$ increases. In particular, for 3-bit models, significant improvements occur even at $k_{chunk}=8$. AWQ’s perplexity decreases from 10.15 to 9.63 for Llama-3 and from 5.96 to 5.53 for Phi-3, while SqueezeLLM’s perplexity drops from 10.49 to 9.93 for Llama-3 and from 5.92 to 5.45 for Phi-3. Meanwhile, the impact on 4-bit models is relatively less pronounced. This is expected, as 4-bit models are already close to full-precision, leaving less room for improvement. The 3.5-bit models follow an intermediate trend.
@@
BBH and MT-Bench. Results on BBH (Figure 14) follow the same trends as the perplexity results. The MT-Bench results (Figure 15), on the other hand, demonstrate some different patterns. In cases where vanilla quantized models without DecDEC ($k=0$) already achieve scores very close to those of the FP16 model—such as all 4-bit cases and the AWQ 3.5-bit model of Phi-3—the scores remain unchanged, oscillating around the baseline score. For the remaining cases, DecDEC significantly improves scores even with a small $k_{chunk}$ (e.g., 8), like in other benchmarks; further increases in $k_{chunk}$, however, do not always yield noticeable improvements. These patterns may be attributed to the coarse-grained rubric of this benchmark, which assigns integer scores ranging from 0 to 10 for each task. This coarse-grained scoring may miss subtle improvements when the potential gain is small.
Figure 16: Comparison against random, static, and exact channel selection. Perplexity is shown on the top rows (lower is better), and recall is shown on the bottom rows (higher is better). The x markers in black correspond to baselines without DecDEC ($k_{chunk}=0$).
Effectiveness of DecDEC’s Channel Selection. Figure 16 compares DecDEC with three variants using different channel selection mechanisms. To assess the benefits of dynamic selection, we include Random (blue line), which selects channels randomly, and Static (yellow line), which statically select channels by Hessian-based ranking with a calibration set (using exact sorting) following prior work [32]. To isolate the effect of DecDEC’s Top-K approximation, we include Exact (red line), which uses true Top-K channels. Alongside perplexity results, we also report the average recall rate of Static and DecDEC relative to Exact.
While Static improves over Random, it underperforms DecDEC. DecDEC achieves lower perplexity than Static while using 4× or even 8× fewer channels (e.g., $k_{\text{chunk}}=32$ or $16$ vs. $128$), highlighting the effectiveness of incorporating input-awareness. Meanwhile, the perplexity gap between DecDEC and Exact is minimal, with almost overlapping curves. These results are explained by the recall rates: while DecDEC achieves around 80% recall relative to Exact, the recall of Static falls significantly short, at around 30% or below.